OpenStudy (anonymous):

Why is Sin(x)/x zero when x=0?

OpenStudy (experimentx):

sin(x)/x is not zero when x=0, it's undefined or inderminate form 0/0

OpenStudy (experimentx):

however, as x->0, sin(x)/x -> 1 we call it limit.

OpenStudy (anonymous):

my bad, i meant to say why is sin(x)/x equal to 1 when x=0

OpenStudy (experimentx):

\( \frac{\sin(x)}{x} \not = 1\) \[ \lim_{x \rightarrow 0} \frac{\sin x}{x}= 1\] http://www.youtube.com/watch?v=Ve99biD1KtA